# IC/DC: Surpassing Heuristic Solvers in Combinatorial Optimization with Diffusion Models

### v1

@@ Abstract
Recent advancements in learning-based combinatorial optimization (CO) methods have shown promising results in solving NP-hard problems without the need for expert-crafted heuristics. However, high performance of these approaches often rely on problem-specific human-expertise-based search after generating candidate solutions, limiting their applicability to commonly solved CO problems such as Travelling Salesman Problem (TSP). In this paper, we present IC/DC, a CO framework that operates without any supervision. IC/DC is specialized in addressing problems involving two distinct sets of items, and it does not need problem-specific search processes to generate valid solutions. IC/DC employs a novel architecture capable of capturing the intricate relationships between items, and thereby enabling effective optimization in challenging CO scenarios. We train our model in a self-supervised way to minimize the cost of the solution while adhering to the problem-specific constraints. IC/DC not only achieves state-of-the-art performance compared to previous learning methods, but also surpasses well-known solvers and heuristic approaches on Asymmetric Traveling Salesman Problem (ATSP).
@@ 1 Introduction
Combinatorial optimization (CO) aims to find the optimal solution that maximizes or minimizes an objective function from a large, discrete set of feasible solutions. This field has been extensively studied due to its broad industrial applications, including logistics, supply chain optimization, job allocation, and more (Zhang et al. 2023). Despite its significance, many CO problems are NP-complete, and developing efficient approximation algorithms is essential.
@@
Traditionally, approximation algorithms for CO have been developed using mathematical programming or hand-crafted heuristics (Miller, Tucker, and Zemlin 1960; Helsgaun 2023). However, the need for problem-specific expertise and the high computational demands of these methods has sparked increasing interest in applying deep learning techniques to CO problems. Early deep learning approaches framed CO problems as sequential decision-making tasks, generating solutions in an auto-regressive manner (Bello et al. 2016; Kool, Van Hoof, and Welling 2018). However, these methods were relatively limited in performance due to their inability to revise previously made decisions.
@@
In contrast, diffusion-based methods generate complete solutions in a single diffusion timestep and then iteratively refine them during the denoising process (Sun and Yang 2023; Min, Bai, and Gomes 2024). This iterative refinement enhances the overall quality of the solutions through corrections and adjustments. By allowing for the revision of earlier decisions, diffusion-based methods overcome the limitations of auto-regressive approaches and avoid the compounding errors typically associated with early decisions.
Despite the impressive performance of diffusion-based methods, previously proposed algorithms had several significant drawbacks, such as the need for costly supervision (Sun and Yang 2023), or use of problem-specific objective that are applicable to other CO problems (Min, Bai, and Gomes 2024). Additionally, due to the challenges in imposing constraint within a diffusion model, previous studies relied on problem-specific search process to extract feasible solutions from the diffusion model’s generations. Designing these problem-specific search requires specialized knowledge and cannot be easily adapted to different CO problems.
In this work, we propose a novel method for training a diffusion model in a self-supervised manner, eliminating the need for costly supervision, problem-specific objectives, or problem-specific search process. We demonstrate our approach on two distinct and challenging CO problems—the parallel machine scheduling problem (PMSP) and the asymmetric travelling salesman problem (ATSP)—which have received less attention (Kwon et al. 2021). Our method not only achieves state-of-the-art performance among deep learning approaches, but also surpasses well-known solvers and heuristic methods on ATSP instances by a significant margin.
@@
Figure 1: The figure illustrates the CO problems we aim to address, and compares various learning-based approaches for CO problems. (Left) We focus on problems involving up to two distinct sets of items, which can be represented as a matrix. In ATSP, rows and columns represent cities (from/to), with each matrix element indicating distance. In PMSP, rows and columns represent machines and jobs, with each element representing processing time. (Center) The diagram highlights how these three approaches differ in generating solutions. Auto-regressive methods generate solutions by iterating through items, while diffusion-based create a heatmap followed by a problem-specific heatmap search process. Our proposed IC/DC approach generates feasible solutions directly, using a training procedure that guides the diffusion model to satisfy the constraints.
@@ 2 Related Works
For widely-studied CO problems like the travelling salesman problem (TSP), several off-the-shelf solvers are available, such as CPLEX (IBM 2022) and OR-tools (Google 2024). These solvers are build on a variety of heuristics, incorporating search methods (Helsgaun 2023), mathematical programming (Arora 1996), and graph algorithms (Christofides 2022). Typically, these approaches rely on problem-specific, handcrafted techniques, which limits their flexibility in adapting to diverse variants encountered in real-world scenarios.
Auto-regressive methods
To overcome this limitation, learning-based solvers have been developed, with early studies primarily focusing on auto-regressive approaches. Bello et al. (2016) were the first to propose solving CO problems using a pointer network trained via reinforcement learning (RL). Kwon et al. (2021) introduced an architecture called MatNet, which builds on the graph attention network (GAT) to encode various types of objects, enabling it to tackle more complex CO problems. While these auto-regressive models offer fast solution generation and can manage intricate CO problems, they are limited by their inability to revise previously made decisions and have been outperformed by methods utilizing diffusion-based methods.
Diffusion-based methods
To the best of our knowledge, two notable works have effectively addressed CO problems using diffusion models, both achieving solution quality comparable to off-the-shelf solvers while significantly reducing generation time. DIFUSCO (Sun and Yang 2023) is a graph neural network (GNN)-based diffusion model trained in a supervised manner to replicate solutions generated by solvers. Similarly, UTSP (Min, Bai, and Gomes 2024) employs a GNN-based diffusion model trained in an unsupervised manner, eliminating the need for costly solution generation from traditional solvers. However, UTSP’s objective is based on the concept of the Hamiltonian cycle, limiting its applicability to TSP.
However, these algorithms fall into the category of heatmap-generating approaches. While CO problems typically impose strict constraints on solutions, such as forming a Hamiltonian cycle, these algorithms are not trained to inherently satisfy those constraints. Instead, they rely on additional heatmap search techniques, such as active search methods (Qiu, Sun, and Yang 2022) or Monte Carlo Tree Search (MCTS) (Silver et al. 2016; Fu, Qiu, and Zha 2021), to produce feasible solutions. This reliance limits their applicability to CO problems with varying constraints on solutions.
@@ 3 Improving Combinatorial Optimization through Diffusion with Constraints
To combine the high-quality solutions of diffusion-based methods with the flexibility of auto-regressive approaches, we propose Improving Combinatorial optimization through Diffusion with Constraints (IC/DC). This approach ensures the feasibility of solutions while training diffusion model in a self-supervised manner, eliminating the need for costly supervision and problem-specific search processes.
@@ 3.1 Problem Definition
We consider a family of CO problems $\mathcal{C}$, which involve two distinct sets of items. Each problem $c\in\mathcal{C}$ is defined by two sets of items $\mathcal{A}$ and $\mathcal{B}$, and matrices that describe the relationships between these two sets of items, as illustrated on the left side of Figure 1. The solution to a CO problem $c$ is represented by a binary matrix $X\in\mathcal{X}=\{0,1\}^{|\mathcal{A}|\times|\mathcal{B}|}$.
@@
$R(X,c)=\begin{cases}\mathrm{score}(X,c)&\text{if $X$ is feasible for $c$}\\ -\infty&\text{otherwise}\end{cases},$
@@ 3.2 Training of IC/DC
We build upon a discrete diffusion model with categorical corruption processes (Austin et al. 2021). We represent the uncorrupted solution that we aim to generate as $X_{0}$, with the corrupted latent variables denoted as $X_{1},...,X_{T}$. We use lowercase $x$ to represent the vectorized forms of $X$, where $x_{t}=\operatorname{vec}(X_{t})\in\{0,1\}^{|\mathcal{A}||\mathcal{B}|}$, and tilded $\tilde{x}$ to denote the one-hot encoded versions, $\tilde{x}_{t}\in\{0,1\}^{|\mathcal{A}||\mathcal{B}|\times 2}$. In line with diffusion model conventions, we use $q(\cdot)$ to denote the data distribution/generative forward process, while $p_{\theta}(\cdot)$ represents the denoising reverse process, which is learned to generate the solutions.
@@
$q(X_{t}|X_{t-1})$ $=\operatorname{Cat}(\operatorname{vec}^{-1}(\tilde{x}_{t-1}Q_{t})),$
$q(X_{t}|X_{0})$ $=\operatorname{Cat}(\operatorname{vec}^{-1}(\tilde{x}{Q}_{1:t})),$
$q(X_{t-1}|X_{t},X_{0})$ $=\operatorname{Cat}\left(\operatorname{vec}^{-1}\left(\frac{\tilde{x}_{t}Q_{t}^{\top}\odot\tilde{x}_{0}Q_{1:t-1}}{\tilde{x}_{0}Q_{1:t}\tilde{x}_{t}^{\top}}\right)\right),$
where $Q_{1:t}=Q_{1}Q_{2}...Q_{t}$, $\odot$ denotes the element-wise multiplication, and $\operatorname{vec}^{-1}$ reshapes the input to the shape $|\mathcal{A}|\times|\mathcal{B}|\times 2$. The matrix $Q_{t}\in[0,1]^{2\times 2}$ is a noise transition matrix that independently applies noise to each element of the solution matrix. We design this noise transition matrix to align with the prior distribution of feasible solutions $\bar{q}$ in the limit (Vignac et al. 2022), such that $\lim_{T\rightarrow\infty}Q_{1:T}z=\bar{q}$ for any vector $z$. This is achieved by defining:
$Q_{t}=\alpha_{t}I+\beta_{t}\mymathbb{1}\bar{q}^{\top},$
@@
We follow the parametrization of Austin et al. (2021), where neural network $f_{\theta}(\cdot)$ is trained to directly predict logits of $X_{0}$ from each $X_{t}$, as follows:
$p_{\theta}(X_{0}|X_{t},c)$ $=\operatorname{Cat}(f_{\theta}(X_{t},t,c)),$
$p_{\theta}(X_{t-1}|X_{t},c)$ $\propto\sum_{X_{0}}q(X_{t-1},X_{t}|X_{0})p_{\theta}(X_{0}|X_{t},c).$
Although this parameterization facilitates the easy computation of diffusion loss when a target dataset is provided, samples from $p_{\theta}$ may not satisfy the constraints since each element of the solution matrix is independently sampled from a Categorical distribution. This limitation required the use of a feasibility-enforcing search process in previous diffusion-based studies on CO (Sun and Yang 2023; Min, Bai, and Gomes 2024). However, this approach requires a search process specifically tailored to each CO problem, and there is no assurance that the search process will preserve the quality of the solution that the diffusion model aims to generate.
Feasibility-enforced generation
To ensure the generation of feasible solutions, we design a process inspired by auto-regressive methods, which we call the feasibility-enforced generation process. In this approach, we sample one element of the solution matrix at a time, ensuring its feasibility based on the previously sampled elements, as follows:
$\widehat{p}_{\theta}(X_{0}|X_{t},c)$ $\propto\prod_{i=1}^{|\mathcal{A}||\mathcal{B}|}\operatorname{Cat}([X_{0}]_{i}|[f_{\theta}(X_{t},t,c)]_{i})\mathbb{I}([X_{0}]_{i}),$
$\mathbb{I}([X_{0}]_{i})$ =$=\begin{cases}1&\text{$[X_{0}]_{i}$ is feasible given $[X_{0}]_{1:i-1}$}\\ 0&\text{otherwise}\end{cases}.$
This approach, similar to the flexibility of auto-regressive methods, allows for the straightforward enforcement of feasibility in the generated samples. However, $\widehat{p}_{\theta}$ cannot be directly utilized as the reverse process because it involves discrete sampling, which prevents the use of the reparameterization trick, thereby making conventional and efficient variational training methods inapplicable. When using $p_{\theta}$ as the reverse process, the connection between $p_{\theta}$ and $\widehat{p}_{\theta}$ becomes weak, leading to a lack of guarantee that samples from $\widehat{p}_{\theta}$ will retain the desired characteristics.
Alternating training
Surrogate targets
$\tilde{q}(X_{0}|c)$ $\propto\exp(R(X_{0},c))[(1-\alpha)q(X_{0})+\alpha p_{\theta}(X_{0}|c)],$
where the mixture is controlled by the hyperparameter $\alpha\in[0,1]$. In the initial stage of training, the solutions generated by reverse process $p_{\theta}(X_{0}|c)$ are not feasible in general. This leads to using more samples from the prior distribution of feasible solutions $q(X_{0})$, and guide the diffusion model to more generate feasible solutions. As training progresses and as $p_{\theta}(X_{0}|c)$ begins to generate feasible solutions, the reward-weighting allows the diffusion model to refine itself by focusing on its high-scoring, feasible generations. Meanwhile, the inclusion of prior distribution of feasible solutions $q(X_{0})$ introduces diversity to the training, counteracting the tendency of $p_{\theta}$ to become too narrow as training progresses.
CLONING step
$\min_{\theta}D_{KL}(\tilde{q}(X_{0}|c)\parallel p_{\theta}(X_{0}|c)),$
which results in a variational bound objective (see Appendix A.2 for detailed derivations), $\mathcal{L}_{\text{VB}}(\theta):=$
$\mathbb{E}_{X_{0}\sim\tilde{q}}\left[\sum_{t=2}^{T}D_{KL}(q(X_{t-1}|X_{t},X_{0})\|p_{\theta}(X_{t-1}|X_{t},c))\right]$
Inspired by recent practices (Austin et al. 2021), we also incorporate the following auxiliary losses:
$\mathcal{L}_{\text{prd}}(\theta)$ $:=\mathbb{E}_{X_{0}\sim\tilde{q}(X_{0}|c),X_{t}\sim q(X_{t}|X_{0})}\left[-\log p_{\theta}(X_{0}|X_{t},c)\right],$
$\mathcal{L}_{\text{cst}}(\theta)$ $:=\mathbb{E}_{X_{0}\sim\tilde{q}(X_{0}|c),X_{t}\sim q(X_{t}|X_{0})}\left[C(p_{\theta}(X_{0}|X_{t},c))\right],$
where $\mathcal{L}_{\text{prd}}$ encourages accurate predictions of the data $X_{0}$ at each time step, and $\mathcal{L}_{\text{cst}}$ discourages infeasible predictions, with $C(\cdot)$ being a differentiable function that approximately measures constraint violations of samples using the Gumbel-softmax trick (see Appendix. A.3 for details on $C$).
In summary, during the CLONING step, we minimize:
IMPROVEMENT step
$\mathcal{L}_{\texttt{IMP}}(\theta)=-\mathbb{E}_{X_{0}\sim\widehat{p}_{\theta}(X_{0}|c)}[R(X_{0},c)].$
Similar to auto-regressive methods, the feasibility-enforced generation process can be viewed as a sequential decision making task, where each element of the solution matrix $X_{0}$ is determined step by step. This perspective allows us to compute the gradient of the above objective using the REINFORCE algorithm (Williams 1992).
where $R^{(i)}=R(X_{0}^{(i)},c)$. This approach is adapted from the baseline estimation method of POMO (Kwon et al. 2020).
Summary
@@
5:     $\theta\leftarrow\theta-\gamma\nabla_{\theta}\mathcal{L}_{\texttt{CLN}}$ according to Eq. (1)
@@
16:     $\theta\leftarrow\theta-\gamma\nabla_{\theta}\mathcal{L}_{\texttt{IMP}}(\theta)$ with $\{(\widehat{X}_{0}^{(i)},c^{(i)})\}_{i=1}^{N}$ according to Eq. (2)

### v4

# Unsupervised Training of Diffusion Models for Feasible Solution Generation in Neural Combinatorial Optimization

@@ Abstract
Recent advancements in neural combinatorial optimization (NCO) methods have shown promising results in generating near-optimal solutions without the need for expert-crafted heuristics. However, high performance of these approaches often rely on problem-specific human-expertise-based search after generating candidate solutions, limiting their applicability to commonly solved CO problems such as Traveling Salesman Problem (TSP). In this paper, we present IC/DC, an unsupervised CO framework that directly trains a diffusion model from scratch. We train our model in a self-supervised way to minimize the cost of the solution while adhering to the problem-specific constraints. IC/DC is specialized in addressing CO problems involving two distinct sets of items, and it does not need problem-specific search processes to generate valid solutions. IC/DC employs a novel architecture capable of capturing the intricate relationships between items, and thereby enabling effective optimization in challenging CO scenarios. IC/DC achieves state-of-the-art performance relative to existing NCO methods on the Parallel Machine Scheduling Problem (PMSP) and Asymmetric Traveling Salesman Problem (ATSP).
@@ 1 Introduction
Combinatorial optimization (CO) aims to find the optimal solution that maximizes or minimizes an objective function from a large, discrete set of feasible solutions. This field has been extensively studied due to its broad industrial applications, including logistics, supply chain optimization, job allocation, and more [32]. Despite its significance, many CO problems are NP-complete, and developing efficient approximation algorithms is essential.
@@
Traditionally, approximation algorithms for CO have been developed using mathematical programming or hand-crafted heuristics [20, 11]. However, the need for problem-specific expertise and the high computational demands of these methods has sparked increasing interest in applying Neural Combinatorial Optimization (NCO), deep learning techniques to CO problems. Early deep learning approaches framed CO problems as sequential decision-making tasks, generating solutions in an autoregressive manner [4, 17]. However, these methods were relatively limited in performance due to their inability to revise previously made decisions.
@@
In contrast, heatmap-based methods construct an initial (possibly infeasible) solution, i.e. heatmap, and iteratively refine the heatmap through corrections and adjustments [26, 21]. By allowing for the revision of earlier decisions, heatmap-based methods overcome the limitations of autoregressive approaches and avoid the compounding errors typically associated with early decisions.
@@
Figure 1: The figure illustrates the CO problems we aim to address, and compares various learning-based approaches for CO problems. (Left) We focus on problems involving up to two distinct sets of items, which can be represented as a matrix. In ATSP, rows and columns represent cities (from/to), with each matrix element indicating distance. In PMSP, rows and columns represent machines and jobs, with each element representing processing time. (Center) The diagram highlights how these three approaches differ in generating solutions. Autoregressive methods generate solutions by iterating through items, while heatmap-based methods create a heatmap followed by a problem-specific heatmap search process. Our proposed IC/DC approach generates feasible solutions directly, using a training procedure that guides the diffusion model to satisfy the constraints.
Despite the impressive performance of heatmap-based methods, previously proposed algorithms had several significant drawbacks, such as the need for costly supervision [26], or use of problem-specific objective that aren’t applicable to other CO problems [21]. More importantly, due to the challenges in imposing constraints on heatmap generation, previous studies relied on problem-specific search process to extract feasible solutions from possibly infeasible heatmaps. Designing these problem-specific search requires specialized knowledge and cannot be easily adapted to different CO problems.
In this work, we propose a novel method for training a feasible solution generator for NCO with a diffusion model in a self-supervised manner, eliminating the need for costly supervision, problem-specific objectives, or problem-specific search process. We demonstrate our approach on two distinct and challenging CO problems: the parallel machine scheduling problem (PMSP) and the asymmetric traveling salesman problem (ATSP), each involving two different classes of items where item features are defined by an asymmetric matrix representing their interrelationships [19]. We show that our method achieves state-of-the-art performance among deep learning approaches.
In summary, the main contributions of this paper are:
  • To the best of our knowledge, this is the first study to introduce a diffusion model for combinatorial optimization (CO) involving two sets of items.
  • We propose IC/DC, a novel method for training a diffusion model in a self-supervised manner, eliminating the need for costly supervision, problem-specific objectives, or problem-specific search process.
  • We empirically demonstrate that our model improves upon other learning-based methods across two distinct CO problems.
@@ 2 Related Works
For widely-studied CO problems like the travelling salesman problem (TSP), several off-the-shelf solvers are available, such as CPLEX [12] and OR-tools [10]. These solvers are build on a variety of heuristics, incorporating search methods [11], mathematical programming [1], and graph algorithms [7]. Typically, these approaches rely on problem-specific, handcrafted techniques, which limits their flexibility in adapting to diverse variants encountered in real-world scenarios.
Autoregressive methods
To overcome this limitation, learning-based solvers have been developed, with early studies primarily focusing on autoregressive approaches. Bello et al. [4] were the first to propose solving CO problems using a pointer network trained via reinforcement learning (RL). Kwon et al. [19] introduced an architecture called MatNet, which builds on the graph attention network (GAT) to encode various types of objects, enabling it to tackle more complex CO problems. While these autoregressive models offer fast solution generation and can manage intricate CO problems, they are limited by their inability to revise previously made decisions. Recent advancements have addressed these limitations by incorporating iterative refinement techniques, mainly utilizing heatmap-based methods and outperformed autoregressive methods in terms of performance.
Heatmap-based methods
To the best of our knowledge, two notable works have effectively addressed CO problems by applying heatmap-based approaches [26, 21]. Both were able to achieve solution quality comparable to off-the-shelf solvers while significantly reducing generation time. DIFUSCO [26] is a graph neural network (GNN)-based diffusion model trained in a supervised manner to replicate solutions generated by solvers. Similarly, UTSP [21] employs a GNN-based model trained in an unsupervised manner, eliminating the need for costly solution generation from traditional solvers. However, UTSP’s objective is based on the concept of the Hamiltonian cycle, limiting its applicability to TSP. While CO problems typically impose strict constraints on solutions, such as forming a Hamiltonian cycle, these algorithms are not trained to inherently satisfy those constraints. Instead, they rely on additional search techniques, such as active search methods [22] or Monte Carlo Tree Search (MCTS) [24, 9], to produce feasible solutions. This reliance limits their applicability to CO problems with varying constraints on solutions.
@@ 3 Improving Combinatorial Optimization through Diffusion with Constraints
To combine the high-quality solutions of heatmap-based methods with the flexibility of autoregressive approaches, we propose Improving Combinatorial optimization through Diffusion with Constraints (IC/DC). This approach ensures the feasibility of solutions while training diffusion model in a self-supervised manner, eliminating the need for costly supervision and problem-specific search processes.
@@ 3.1 Problem Definition
We consider a family of CO problems $\mathcal{C}$, which involve two distinct sets of items. Each problem $c\in\mathcal{C}$ is defined by two sets of items $\mathcal{A}$ and $\mathcal{B}$, and matrices that describe the relationships between these two sets of items, as illustrated on the left side of Fig. 1. The solution to a CO problem $c$ is represented by a binary matrix $X\in\mathcal{X}=\{0,1\}^{|\mathcal{A}|\times|\mathcal{B}|}$.
@@
$R(X,c)=\begin{cases}\mathrm{score}(X,c)&\text{if $X$ is feasible for $c$}\\ -\infty&\text{otherwise}\end{cases},$ (1)
@@ 3.2 Training of IC/DC
We build upon a discrete diffusion model with categorical corruption processes [2]. We represent the uncorrupted solution that we aim to generate as $X_{0}$, with the corrupted latent variables denoted as $X_{1},...,X_{T}$. We use lowercase $x$ to represent the vectorized forms of $X$, where $x_{t}=\operatorname{vec}(X_{t})\in\{0,1\}^{|\mathcal{A}||\mathcal{B}|}$, and tilded $\tilde{x}$ to denote the one-hot encoded versions, $\tilde{x}_{t}\in\{0,1\}^{|\mathcal{A}||\mathcal{B}|\times 2}$. In line with diffusion model conventions, we use $q(\cdot)$ to denote the data distribution/generative forward process, while $p_{\theta}(\cdot)$ represents the denoising reverse process, which is learned to generate the solutions.
@@
$q(X_{t}|X_{t-1})$ $=\operatorname{Cat}(\operatorname{vec}^{-1}(\tilde{x}_{t-1}Q_{t})),$ (2)
$q(X_{t}|X_{0})$ $=\operatorname{Cat}(\operatorname{vec}^{-1}(\tilde{x}{Q}_{1:t})),$ (3)
$q(X_{t-1}|X_{t},X_{0})$ $=\operatorname{Cat}\left(\operatorname{vec}^{-1}\left(\frac{\tilde{x}_{t}Q_{t}^{\top}\odot\tilde{x}_{0}Q_{1:t-1}}{\tilde{x}_{0}Q_{1:t}\tilde{x}_{t}^{\top}}\right)\right),$ (4)
where $Q_{1:t}=Q_{1}Q_{2}...Q_{t}$, $\odot$ denotes the element-wise multiplication, and $\operatorname{vec}^{-1}$ reshapes the input to the shape $|\mathcal{A}|\times|\mathcal{B}|\times 2$. The matrix $Q_{t}\in[0,1]^{2\times 2}$ is a noise transition matrix that independently applies noise to each element of the solution matrix. We design this noise transition matrix to align with the prior distribution of feasible solutions $\bar{q}$ in the limit [29], such that $\lim_{T\rightarrow\infty}Q_{1:T}z=\bar{q}$ for any vector $z$. This is achieved by defining:
$Q_{t}=\alpha_{t}I+\beta_{t}\mymathbb{1}\bar{q}^{\top},$ (5)
@@
We follow the parametrization of Austin et al. [2], where neural network $f_{\theta}(\cdot)$ is trained to directly predict logits of $X_{0}$ from each $X_{t}$, as follows:
$p_{\theta}(X_{0}|X_{t},c)$ $=\operatorname{Cat}(f_{\theta}(X_{t},t,c)),$ (6)
$p_{\theta}(X_{t-1}|X_{t},c)$ $\propto\sum_{X_{0}}q(X_{t-1},X_{t}|X_{0})p_{\theta}(X_{0}|X_{t},c).$ (7)
@@
5:     $\theta\leftarrow\theta-\gamma\nabla_{\theta}\mathcal{L}_{\texttt{CLN}}$ according to Eq. 15
@@
16:     $\theta\leftarrow\theta-\gamma\nabla_{\theta}\mathcal{L}_{\texttt{IMP}}(\theta)$ with $\{(\widehat{X}_{0}^{(i)},c^{(i)})\}_{i=1}^{N}$ according to Eq. 17
@@
Although this parameterization facilitates the easy computation of diffusion loss when a target dataset is provided, samples from $p_{\theta}$ may not satisfy the constraints since each element of the solution matrix is independently sampled from a Categorical distribution. This limitation required the use of a feasibility-enforcing search process in previous heatmap-based studies on CO [26, 21]. However, this approach requires a search process specifically tailored to each CO problem, and there is no assurance that the search process will preserve the quality of the solution that the diffusion model aims to generate.
Feasibility-enforced generation
To ensure the generation of feasible solutions, we design a process inspired by autoregressive methods, which we call the feasibility-enforced generation process. In this approach, we sample one element of the solution matrix at a time, ensuring its feasibility based on the previously sampled elements, as follows:
$\widehat{p}_{\theta}(X_{0}|X_{t},c)$ $\propto\prod_{i=1}^{|\mathcal{A}||\mathcal{B}|}\operatorname{Cat}([X_{0}]_{i}|[f_{\theta}(X_{t},t,c)]_{i})\mathbb{I}([X_{0}]_{i}),$ (8)
$\mathbb{I}([X_{0}]_{i})$ =$=\begin{cases}1&\text{$[X_{0}]_{i}$ is feasible given $[X_{0}]_{1:i-1}$}\\ 0&\text{otherwise}\end{cases}.$ (9)
This approach, similar to the flexibility of autoregressive methods, allows for the straightforward enforcement of feasibility in the generated samples. However, $\widehat{p}_{\theta}$ cannot be directly utilized as the reverse process because it involves discrete sampling, which prevents the use of the reparameterization trick, thereby making conventional and efficient variational training methods inapplicable. When using $p_{\theta}$ as the reverse process, the connection between $p_{\theta}$ and $\widehat{p}_{\theta}$ becomes weak, leading to a lack of guarantee that samples from $\widehat{p}_{\theta}$ will retain the desired characteristics.
Alternating training
Surrogate targets
$\tilde{q}(X_{0}|c)$ $\propto\exp(R(X_{0},c))[(1-\alpha)q(X_{0})+\alpha p_{\theta}(X_{0}|c)],$ (10)
where the mixture is controlled by the hyperparameter $\alpha\in[0,1]$. In the initial stage of training, the solutions generated by reverse process $p_{\theta}(X_{0}|c)$ are not feasible in general. This leads to using more samples from the prior distribution of feasible solutions $q(X_{0})$, and guide the diffusion model to generate feasible solutions. As training progresses and as $p_{\theta}(X_{0}|c)$ begins to generate feasible solutions, the reward-weighting allows the diffusion model to refine itself by focusing on its high-scoring, feasible generations. Meanwhile, the inclusion of prior distribution of feasible solutions $q(X_{0})$ introduces diversity to the training, counteracting the tendency of $p_{\theta}$ to become too narrow as training progresses.
@@
CLONING step
$\min_{\theta}D_{KL}(\tilde{q}(X_{0}|c)\parallel p_{\theta}(X_{0}|c)),$ (11)
which results in a variational bound objective (see Sec. A.2 for detailed derivations), $\mathcal{L}_{\text{VB}}(\theta):=$
$\mathbb{E}_{X_{0}\sim\tilde{q}}\left[\sum_{t=2}^{T}D_{KL}(q(X_{t-1}|X_{t},X_{0})\|p_{\theta}(X_{t-1}|X_{t},c))\right]$ (12)
Inspired by recent practices [2], we also incorporate the following auxiliary losses:
$\mathcal{L}_{\text{prd}}(\theta)$ $:=\mathbb{E}_{X_{0}\sim\tilde{q}(X_{0}|c),X_{t}\sim q(X_{t}|X_{0})}\left[-\log p_{\theta}(X_{0}|X_{t},c)\right],$ (13)
$\mathcal{L}_{\text{cst}}(\theta)$ $:=\mathbb{E}_{X_{0}\sim\tilde{q}(X_{0}|c),X_{t}\sim q(X_{t}|X_{0})}\left[C(p_{\theta}(X_{0}|X_{t},c))\right],$ (14)
where $\mathcal{L}_{\text{prd}}$ encourages accurate predictions of the data $X_{0}$ at each time step, and $\mathcal{L}_{\text{cst}}$ discourages infeasible predictions, with $C(\cdot)$ being a differentiable function that approximately measures constraint violations of samples using the Gumbel-softmax trick (see Sec. A.3 for details on $C$).
In summary, during the CLONING step, we minimize:
IMPROVEMENT step
$\mathcal{L}_{\texttt{IMP}}(\theta)=-\mathbb{E}_{X_{0}\sim\widehat{p}_{\theta}(X_{0}|c)}[R(X_{0},c)].$ (16)
Similar to autoregressive methods, the feasibility-enforced generation process can be viewed as a sequential decision making task, where each element of the solution matrix $X_{0}$ is determined step by step. This perspective allows us to compute the gradient of the above objective using the REINFORCE algorithm [30].
where $R^{(i)}=R(X_{0}^{(i)},c)$. This approach is adapted from the baseline estimation method of POMO [18].
Summary